Accelerator Update for February 13, 2009 - February 16, 2009 |
The Integrated Luminosity for 2/9/09 to 2/16/09 was ~23.74 Pb-1. Friday Feb 13 The day shift began with the accelerator complex preparing for a MCR power outage and with many machines and experiments in access. The High Voltage Coordinator turned off the power to the Cross Gallery, MCR, and computer rooms at 8:30 AM to connect the new transformer. The High Voltage Coordinator turned the power on to the Cross Gallery, MCR, and the computer rooms around 10:50 AM. Operations reported at 11:54 AM that the TCLK was back and the CAMAC crates were restored. Controls experts reported at 12:27 PM that everything looked “good.” The D-Zero (D0) experiment came out of access at 12:50 PM. The Main Injector (MI) came out of access around 1:17 PM. At 2:03 PM, Operations reported that they were having problems closing the MI-20 VCB. A power supply expert investigated and fixed the problem. The Collider Detector at Fermilab (CDF) came out of access at 3:11 PM. At 3:45 PM, a Tevatron (TeV) expert reported that the dump switches at sectors A3 and E2 needed work. Operations established Linac beam at 3:48 PM. At 4:07 PM, a Linac quadrupole power supply (QPS102) tripped off. A Linac expert investigated and brought it back online. Operations established beam to Booster and the MiniBooNE experiment at 4:27 PM. Operations reported that the Accumulator had retained a stack of 135.7E10 antiprotons through the power outage. The Antiproton Source resumed stacking at 4:43 PM. At 5:08 PM, experts reported that the TeV sectors A3 and E2 dump switches failed their dump test. A Pbar expert began tuning the Accumulator and Debuncher at 5:17 PM. At 5:26 PM, a Linac RF station (LRF1) tripped off. Experts had the station back up by 5:54 PM. Operations reported at 6:08 PM that the Recycler’s Electron Cooling system was up and running. At 7:09 PM, Operations reported a problem with a MI power supply (HV703). Operations reported at 7:26 PM that all the TeV dump switches were working. At 7:42 PM, the TeV sector C3 dump switch failed its test. An expert investigated. Operations reported at 7:53 PM that HV703 was not decoding the TCLK properly. This was corrected. A TeV expert reported at 8:04 PM that all the rest of the dump switch controllers would have to be updated so that they matched the two at sectors A3 and E2. They will have to replace all the series SCRs around the ring. Operations reported at 8:19 PM that they had stashed antiprotons twice. At 9:01 PM, LRF3 tripped off. It reset. At 9:07 PM, Operations reported that TeV experts have discovered more dump switches that were damaged by the power outage last Tuesday night. They will begin replacing the SCRs on Saturday morning. Operations reported at midnight that the Recycler had stashed antiprotons four times during the evening shift. Saturday Feb 14 The midnight shift began with Operations monitoring the stash (198.3E10), with Pbar stacking (24.2E10), and with NuMI and MiniBooNE taking beam. At 5:03 AM, LRF3 tripped off. It reset. Operations reported at 8 AM that the Recycler had stashed nine times during the midnight shift. Cryo system experts from CHL reported at 10:34 AM that the TeV sector C4 wet engine needed some maintenance. It took them about 20 minutes. Operations reported at 2:10 PM that the TeV had successfully completed a hipot. TeV experts continued to work on the dump switches. At 3:27 PM, a Booster kicker (MKS10) tripped off due to an internal cooling fan failure. Operations bypassed the kicker and called in an expert. A Booster expert reported at 5:02 PM that he had fixed the fan for MKS10 and brought the kicker back online. Operations turned off the TeV at 6:37 PM so that experts could repair a cable. Operations reported at 8:02 PM that the Recycler had stashed two times during the last 12 hours. Operations reported at 8:55 PM that they had successfully completed a TeV dump test at 150 GeV. TeV experts began conducting thermal scans of all the knife switches. Operations ramped the TeV at 10:30 PM. Operations put the TeV through a dry squeeze at 10:44 PM. At 11:28 PM, the TeV suffered a quench at sector F1 during an injection of protons. Sunday Feb 15 The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. A TeV expert reported at 1:14 AM that he believed a controls card had corrupted data and this was what caused the quench. Operators changed the card. Operations ramped the TeV at 2:52 AM. Operations put the TeV through a dry squeeze at 3:05 AM. Operations put the TeV through a wet squeeze and then conducted an orbit check at 3:56 AM. At 4:30 AM, Operations reported that they could not turn on the Tevatron Electron Lens (TEL2). Operations called in an expert who came in and repaired it. The TEL2 expert reported at 6:56 AM that its modulator was broken. Operations called in another expert. Operations put the TeV through a wet squeeze at 8:40 AM. At 9:31 AM, Operations reported that the TeV sequencer failed while checking on BLM states. Operations contacted an expert who said not to worry about it. Operations put the TeV into shot setup at 9:45 AM. At 11:04 AM, Operations reported that earlier a Linac Klystron RF station (KRF6) tripped off on a LCW flow fault. It eventually reset. Operations put the TeV through a wet squeeze at 1:08 PM. Operations put the TeV into a new shot setup at 1:31 PM. At 1:49 PM, the P1 transfer line suffered a vacuum burst. Operations disabled beam and closed vacuum vacuum. About 1.5 hours later the vacuum got good enough to run beam and resume the shot setup. Operations established store 6809 at 4:46 PM with an initial luminosity of 324.2E30. Operations reported at 8:06 PM that the Recycler had stashed five times during the last 12 hours. Operations turned on the Switchyard and Meson power supplies at 11:28 PM in preparation to send beam to the M01 dump on Monday morning. Monday Feb 16 The midnight shift began with Operations monitoring store 6809 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. At 12:22 AM, Operations reported that the 1.7GHz TeV tune Schottky was not operating. Operations will contact experts later in the morning. Operations terminated store 6809 sometime around 5 AM. Operations put the TeV into shot setup at 5:20 AM. Operations established store 6810 at 7:01 AM with an initial luminosity of 323.48E30. Operations reported at 7:58 AM that the Recycler had stashed nine times during the last 12 hours. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 02/16/2009 email Fermilab |
| Security, Privacy, Legal | |